# Improved multifidelity Monte Carlo estimators based on normalizing flows and dimensionality reduction techniques

### v1

@@ 4.1 Analytic functions
    Inspired by [10], we consider the following analytic functions as high-fidelity and low-fidelity models
@@ 4.2 Reaction-diffusion equation
    and we set the final time $T=4$ and the size of the domain $L=1$. The equations are solved employing the finite volume method for space discretization and the fourth order Runge–Kutta method for time integration. The high-fidelity model is obtained by setting 64 finite volume cells in each direction and running the simulation for 400 time steps, while for the low-fidelity model we use 16 finite volume cells in each direction and run for 100 time steps, which yields a cost ratio of $w=0.1$. Moreover, in the low-fidelity model we replace the diffusion coefficients $D_{u}$ and $D_{v}$ by their average value, i.e., $\bar{D}=(D_{u}+D_{v})/2$. The input parameters for the high-fidelity model are the diffusion coefficients $D_{u},D_{v}$ and the reaction coefficient $k$, while for the low-fidelity model we have $\bar{D}$ and $k$, and therefore we are in the framework where the two models have a dissimilar parameterization. In Fig. 10 we plot the solutions $u$ and $v$ of equations (4.5) for both the high-fidelity and low-fidelity models, setting the parameters $D_{u}=10^{-3},D_{v}=5\cdot 10^{-3},k=10^{-3}$, and $\bar{D}=3\cdot 10^{-3}$. The input probability distribution for the forward uncertainty propagation study is such that $k$ is independent of $D_{u}$ and $D_{v}$, $k\sim\mathcal{U}([0.5\cdot 10^{-3},1.5\cdot 10^{-3}])$, and $D_{u},D_{v}$ are uniformly distributed in the triangle with vertices $\{(0.25\cdot 10^{-3},4\cdot 10^{-3}),(1.75\cdot 10^{-3},5\cdot 10^{-3}),(1\cdot 10^{-3},6\cdot 10^{-3})\}$. The distribution of $\bar{D}$ is obtained accordingly.
@@ 4.3 Cardiovascular simulation
    5 Conclusion
    In this work we proposed two different methodologies to improve multifidelity estimators for uncertainty propagation. In particular, we achieved variance reduction of the standard multifidelity Monte Carlo estimator by modifying the low-fidelity model in order to increase the correlation with the high-fidelity model. Our approaches rely on a shared space where the models vary the most and the parameters are distributed according to a standard Gaussian. We constructed the shared space through either linear or nonlinear dimensionality reduction techniques, namely active subspaces and autoencoders. We demonstrated by means of numerical experiments that, given sufficient data, the latter are able to find nonlinear transformations which allow us to further decrease the variance of the estimator with respect to linear transformations. Moreover, we employed normalizing flows to map different probability distributions into the same distribution, i.e., a standard Gaussian, and therefore generate a shared space. Our techniques not only permit to get an estimator with reduced variance, but also increase the range of applicability of multifidelity Monte Carlo estimators. In particular, we allow for models with dissimilar parameterization, meaning that the number and type of input parameters between the high-fidelity and low-fidelity models and their distributions can be different. This implies that these approaches can also be applied to models which measure different quantities of interest which are not directly related, as long as the high-fidelity and the modified low-fidelity models are well correlated, as we showed in the challenging numerical examples involving cardiovascular simulations. As a future development we are interested in studying a method to automatically find the optimal reduced dimension of the shared space, which is fundamental for applications with high-dimensional input parameters.
@@ Acknowledgements
    This work is supported by NSF CAREER award #1942662 (DES), NSF CDS&E award #2104831 (DES), NSF award #2105345 (ALM), and NIH grant #5R01HL141712 (ALM, KM). This work used computational resources from the Stanford Research Computing Center (SRCC). This article has been authored by an employee of National Technology & Engineering Solutions of Sandia, LLC under Contract No. DE-NA0003525 with the U.S. Department of Energy (DOE). The employee owns all right, title and interest in and to the article and is solely responsible for its contents. The United States Government retains and the publisher, by accepting the article for publication, acknowledges that the United States Government retains a non-exclusive, paid-up, irrevocable, world-wide license to publish or reproduce the published form of this article or allow others to do so, for United States Government purposes. The DOE will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan https://www.energy.gov/downloads/doe-public-access-plan.

### v2

@@ 4.1 Analytic functions
    Inspired by [12], we consider the following functions as high-fidelity and low-fidelity models
@@ 4.2 Reaction-diffusion equation
    and we set the final time $T=4$ and the size of the domain $L=1$. The equations are solved employing the finite volume method for space discretization and the fourth order Runge–Kutta method for time integration. The high-fidelity model is obtained by setting 64 finite volume cells in each direction and running the simulation for 400 time steps, while for the low-fidelity model we use 16 finite volume cells in each direction and run for 100 time steps, which yields a cost ratio of $w=0.1$. Moreover, in the low-fidelity model we replace the diffusion coefficients $D_{u}$ and $D_{v}$ by their average value, i.e., $\bar{D}=(D_{u}+D_{v})/2$. The input parameters for the high-fidelity model are the diffusion coefficients $D_{u},D_{v}$ and the reaction coefficient $k$, while for the low-fidelity model we have $\bar{D}$ and $k$, and therefore we are in the framework where the two models have a dissimilar parameterization. In Fig. 10 we plot the solutions $u$ and $v$ of equations (4.5) for both the high-fidelity and low-fidelity models, setting the parameters $D_{u}=10^{-3},D_{v}=5\cdot 10^{-3},k=10^{-3}$, and $\bar{D}=3\cdot 10^{-3}$. The input probability distribution for the forward uncertainty propagation study is such that $k$ is independent of $D_{u}$ and $D_{v}$, $k\sim\mathcal{U}([0.5\cdot 10^{-3},1.5\cdot 10^{-3}])$, and $D_{u},D_{v}$ are uniformly distributed in the triangle with vertices $\{(0.25\cdot 10^{-3},4\cdot 10^{-3}),(1.75\cdot 10^{-3},5\cdot 10^{-3}),(1\cdot 10^{-3},6\cdot 10^{-3})\}$. We remark that this is an example of correlated random inputs, since $D_{u}$ and $D_{v}$ in the high-fidelity model are uniformly distributed on a triangular support, leading to correlated parameters. The distribution of $\bar{D}$ is then obtained accordingly. One of the challenges when working with dependent inputs relates to the difficulty of enforcing that the new low-fidelity samples generated by the pipeline are defined over the same support as the original low-fidelity inputs. To overcome this problem, we assume the low-fidelity model to be defined even outside the original support of input parameters, and we are currently investigating how to lift such assumption.
@@ 4.3 Cardiovascular simulation
    5 Conclusions
    In this work we proposed two different methodologies to improve multifidelity estimators for uncertainty propagation. In particular, we achieved variance reduction of the standard multifidelity Monte Carlo estimator by modifying the low-fidelity model in order to increase the correlation with the high-fidelity model. Our approaches rely on a shared space where the models vary the most and the parameters are distributed according to a standard Gaussian. We constructed the shared space through either linear or nonlinear dimensionality reduction techniques, namely active subspaces and autoencoders. We demonstrated by means of numerical experiments that, given sufficient data, the latter are able to find nonlinear transformations which allow us to further decrease the variance of the estimator with respect to linear transformations. Moreover, we employed normalizing flows to map different probability distributions into the same distribution, i.e., a standard Gaussian, and therefore generate a shared space. Our techniques not only permit getting an estimator with reduced variance, but also increase the range of applicability of multifidelity estimators. In particular, we allow for models with dissimilar parameterization, meaning that the number and type of input parameters between the high-fidelity and low-fidelity models and their distributions can be different. This implies that these approaches can also be applied to models which measure different quantities of interest which are not directly related, as long as the high-fidelity and the modified low-fidelity models are well correlated, as we showed in the challenging numerical examples involving cardiovascular simulations. A limitation of our approach is that a large amount of data might be necessary to train the surrogate models, find the lower-dimensional subspaces, and build the normalizing flows. However, as we showed in Figs. 6 and 8 where we varied the number of data points, even if the best lower-dimensional manifolds are not correctly identified, we still have an improvement in terms of variance with respect to standard multifidelity Monte Carlo estimators. In other words, through the examples in the paper, we numerically show that the number of samples needed to build an accurate surrogate is typically much larger than those needed to build a surrogate for the sole purpose of identifying a shared space leading to a smaller variance than multifidelity Monte Carlo. The sensitivity analysis with respect to the number of samples also shows that, as expected, the variance of the resulting estimators decreases with a larger amount of data. Moreover, the construction of an accurate surrogate model could be lifted in the linear dimension reduction case by adopting strategies like adaptive basis (see, e.g., [45]) or it could be directly learned on the latent space together with the autoencoder. We leave this latter approach, which we expect to require a significantly smaller amount of data due to the reduced dimension, for future work. Another interesting extension of the current method is to train the autoencoders of both the high-fidelity and low-fidelity models together, and include a term in the loss function that maximizes the resulting correlation and therefore improves the variance of the estimators. Finally, we are also interested in studying a method to automatically find the optimal reduced dimension of the shared space, which is fundamental for applications with high-dimensional input parameters.
@@ Acknowledgements
    This work is supported by NSF CAREER award #1942662 (DES), NSF CDS&E award #2104831 (DES), NSF award #2105345 (ALM), and NIH grant #5R01HL141712 (ALM, KM). This work used computational resources from the Stanford Research Computing Center (SRCC). Sandia National Laboratories is a multi-mission laboratory managed and operated by National Technology & Engineering Solutions of Sandia, LLC (NTESS), a wholly owned subsidiary of Honeywell International Inc., for the U.S. Department of Energy’s National Nuclear Security Administration (DOE/NNSA) under contract DE-NA0003525. This written work is authored by an employee of NTESS. The employee, not NTESS, owns the right, title and interest in and to the written work and is responsible for its contents. Any subjective views or opinions that might be expressed in the written work do not necessarily represent the views of the U.S. Government. The publisher acknowledges that the U.S. Government retains a non-exclusive, paid-up, irrevocable, world-wide license to publish or reproduce the published form of this written work or allow others to do so, for U.S. Government purposes. The DOE will provide public access to results of federally sponsored research in accordance with the DOE Public Access Plan. The authors thank Boris Kramer for insightful suggestions about Section 3.3.